Pure water has h+ ion concentration as 10^-6 M. The value of Kw at this temperature is

Grade:12

2 Answers

Sudhir Bishnoi
6 years ago
For pure water [H+]=[OH-]
Here [H+]=10-6=[OH-]
Since Kw=[H+][OH-]
           Kw=10-6*10-6=10-12
Hence for water Kw is 10-12
Chhavi Jain
6 years ago
Ionic product of water is given as :
K= [H+] [OH-]
In case of pure water : [H+] = [OH-]
K= 10-6  * 10-6
      = 10-12